OpenStudy (anonymous):

Harriet is comparing the costs of calling cards. To use card A, the cost is $1.00 to connect and then $0.04 per minute. To use card B, it costs $0.65 to connect and then $0.06 per minute. For what number of minutes does it cost the same amount to use each card for a single call? (Enter only the number.)

OpenStudy (ybarrap):

Card A costs A=1 + .04m, where m is minutes Card B costs B=.65 + .06m, where m is minutes At some point, A=B. Just equate these two equations and solve for m to see when this condition happens:

OpenStudy (ybarrap):

Make sense?

OpenStudy (anonymous):

Um... a little @ybarrap

OpenStudy (ybarrap):



OpenStudy (ybarrap):

$$ A=B\\ \implies 1 + .04m=.65 + .06m\\ \text{ So, }\\ 1-0.65=.06m-0.04m=m(0.06-0.04)\\ \implies m=\cfrac{1-0.65}{0.06-0.04} $$ At this value of \(m\), \(A=B\).